OpenStudy (anonymous):

Factor the polyomial completely using any method y^4-81 Medal given for answer

OpenStudy (anonymous):

(y^2)^2 - (9)^2 = (Y^2 - 9) (y^2+ 9)

OpenStudy (anonymous):

further

OpenStudy (anonymous):

Thanks medal given (:

OpenStudy (anonymous):

(y^2)^2 - (9)^2 = (Y^2 - 9) (y^2+ 9) = [(y)^2 - (3)^2] (y^2 + 9) = (y-3)(y+3)(y^2 + 9)

OpenStudy (anonymous):

pls see my final ans above......☺ thanks for the medal....